A motor rotating at 500 rpm is switched off. It comes to rest in 10 seconds. Find the angular retardation and the number of revolutions made by it before stopping.

Grade:11

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

Given , wo= 500 rpm = 500 /60 rotations per second = 25/3 rps
time = 10 seconds

So, w = 0 = wo+\alphat,
\alpha= – 25/ 30 rotations / s2= -5/6 rotations / s2

Now,\Theta= wo t+ 0.5\alphat2

So,\Theta= (25/3) 10 + 0.5 (-5/6) 102= 125/3 rotations.

thanks